i need a word problem that is created that gives hints about what the angle measurement is. it is for a triangle so the sum of all of them is 180 degrees. A EXAMPLE THAT WAS ON MY HOMEWORK WAS:
angle 2 is three more than the first and angle three is one more than twice the
first. What are the measurements of the three angles.... i need something like this to get the word problem please email me if you know a good on im confused
Thanks

Answer by checkley79(3341)

ANGLE 2(Y)=ANGLE 1(X) +3 DEGREE.
ANGLE 3(Z)= 2 TIMES THE (X)+ 1
X+Y+Z=180
X+(X+3)+(2X+1)=180
4X+4=180
4X=180-4
4X=176
X=176/4
X=44 THE FIRST ANGLE.
Y=44+3=47 FOR ANGLE 2.
2*44+3=88+1=91 FOR THE THIRD ANGLE.
PROOF:
44+47+89=180
180=180